The difference of Ron's age and Joan's ages is 8 years. The sum of twice Ron's age and 3 times Joan's age is 66. How old is each.
Answer by mananth(16946)

The difference of Ron's age and Joan's ages is 8 years. The sum of twice Ron's age and 3 times Joan's age is 66. How old is each.
..
Ron's age =r
Joan's age =j
..
r-j=8...................1
2r+3j=66................2
..
multiply 1 by 3
3r-3j=24
add this to (2)
2r+3r+3j-3j=24+66
5r=90
/5
r=14 -----ron's age
Joan's age =6
...
m.ananth@hotmail.ca